OpenStudy (anonymous):

Solve. 7^(x – 5) = 7^(4x + 7)

OpenStudy (anonymous):

x-5=4x+7===> 3x=-12 ==> x=-12/3 =-4

OpenStudy (anonymous):

simplifies to x-5 = 4x +7

OpenStudy (anonymous):

thanx